OpenStudy (anonymous):

0.3x-0.2y=4 0.4x+0.3y= 73/19

OpenStudy (anonymous):

{3/10 x - 2/10 y = 4, 4/10 x + 3/10 y = 73/19}\[\left\{x\to \frac{220}{19},y\to -\frac{50}{19}\right\} \]

OpenStudy (saifoo.khan):

x = 11.57 y = -2.63

OpenStudy (angela210793):

0.3x-0.2y=4(*3) 0.9x-0.6y=12 1.7x=146/19+12 0.4x+0.3y= 73/19(*2) 0.8x+0.6y= 146/19(add both sides) 1.7x=374/19 x=11.5 .3*11.5-0.2y=4

OpenStudy (anonymous):

The following Mathematica expression replaces the x and y solutions into both equations. If the LHS is equal to the RHS, then Mathematica returns "True". Some of the original fractions have been simplified.\[\left\{\frac{3 x}{10}-\frac{y}{5}=4,\frac{2 x}{5}+\frac{3 y}{10}=\frac{73}{19}\right\}\text{/.}\left\{x\to \frac{220}{19},y\to -\frac{50}{19}\right\} \to \{\text{True},\text{True}\} \]The original coefficients were rationalized. Why would one do that? Well the answers are exact.